# Machine-learning for photoplethysmography analysis: Benchmarking feature, image, and signal-based approaches

### v1

@@ Abstract
Photoplethysmography (PPG) is a widely used non-invasive physiological sensing technique, suitable for various clinical applications. Such clinical applications are increasingly supported by machine learning methods, raising the question of the most appropriate input representation and model choice. Comprehensive comparisons, in particular across different input representations, are scarce. We address this gap in the research landscape by a comprehensive benchmarking study covering three kinds of input representations, interpretable features, image representations and raw waveforms, across prototypical regression and classification use cases: blood pressure and atrial fibrillation prediction. In both cases, the best results are achieved by deep neural networks operating on raw time series as input representations. Within this model class, best results are achieved by modern convolutional neural networks (CNNs). but depending on the task setup, shallow CNNs are often also very competitive. We envision that these results will be insightful for researchers to guide their choice on machine learning tasks for PPG data, even beyond the use cases presented in this work.
@@ 1 Introduction
PPG is one of the most commonly used wearable sensing techniques. It consists of projecting light onto the skin and measuring the amount of light that is reflected back or transmitted through the underlying tissues. Its simplicity, non-invasive nature, and ability to deliver multiple physiological parameters make it particularly attractive [1, 2]. As a result, PPG has been integrated into a range of clinical devices, such as pulse oximeters, as well as consumer wearable devices including smartwatches.
The PPG signal measures the fluctuations in blood volume in the skin’s microvascular bed which occur with each heartbeat. Figure 1 shows an exemplary PPG signal: the shape of the pulse wave contains information relating to the heart and vasculature, including blood pressure; and the inter-beat intervals are related to heart rhythm [3]. The time delay between the electrical activation of the heart and the arrival of the corresponding pulse wave at a peripheral site where PPG is measured has been used to predict blood pressure There are generally two approaches to analysing PPG signals (1): (i) using signal processing to extract features relating to pulse wave shape or inter-beat-intervals; and (ii) using deep learning techniques to analyse PPG signals or their image-based representations [4]. Over the past few years, deep learning techniques have become widely used [5, 6].
@@
In this work, we investigate two widely considered clinical applications for PPG analysis: atrial fibrillation (AF) classification as a prototypical classification task and cuffless blood pressure (BP) estimation as a prototypical regression task. AF is the most common sustained cardiac arrhythmia and confers a five-fold increase in stroke risk [8]. AF is characterised by irregular and often very rapid heart rhythm. PPG provides an attractive approach to identifying AF because it is widely used in consumer wearables, and because it can detect each heart beat, it can provide measures of the heart rhythm. In the case of AF, the time intervals between heart beats are irregular. BP is one of the most widely used physiological measurements. It is a key marker of cardiovascular health; a valuable predictor of cardiovascular events; and is essential for the selection and monitoring of antihypertensive (BP lowering) treatments [9]. PPG-based blood pressure estimation provides a potential approach to monitor BP unobtrusively in daily life. The rationale behind selecting two prototypical, but very different use cases is to identify general patterns that could guide practitioners in the field even beyond the two investigated use cases.
While many prediction models have been put forward for the two considered prediction tasks at hand, benchmarking results are typically presented within a set of prediction models operating on a single kind of input modality or a comparison is carried out against previously reported results from the literature. However, the latter rely on matching the experimental setup as closely as possible, where deviations from this setup severely limit the comparability of the results. The current lack of like-to-like comparisons is an important gap in the research landscape, that we aim to address with this submission.
In this work, we address the following research questions: How do state-of-the-art machine learning models operating on different input representations compare? Are there universal patterns in terms of best-performing input representations or model architectures across different prototypical classification and regression use cases? As our main technical contribution, we put forward a like-to-like benchmarking of a comprehensive set of state-of-the-art algorithms covering three kinds of input representations on two large-scale datasets for a prototypical classification task (AF classification) and two different variants of a prototypical regression (BP regression) task.
@@ 2 Related Work
Machine Learning Approaches. Typically, all approaches for clinical prediction tasks based on PPG data rely on a combination of signal processing and machine learning, where the precise focus varies considerably across different approaches. At one end of the spectrum, approaches relying on clinically interpretable features focus heavily on signal processing to extract meaningful features and typically use comparably simple classification/regression models to perform the prediction. At the other end of the spectrum, deep learning methods using raw time series as input with as little signal processing as possible rely on complex prediction models to extract and process meaningful features by themselves. In between, there are classifiers based on image-representations, that leverage signal processing tools to turn raw waveforms into image representations and then most commonly also rely on deep learning models to perform the prediction.
@@
Time Series Models. Recent advancements in deep learning have significantly impacted healthcare by enabling complicated analysis of raw physiological data. These models are particularly effective in estimating BP and detecting AF, among other applications. The key advantage of deep learning is its ability to recognize complicated patterns directly from raw data such as electrocardiogram (ECG) and PPG signals, eliminating the need for extensive manual feature development [10, 11]. They provide an effective means to learn the complex, nonlinear underlying relationship between PPG signals and various physiological parameters, without the need to define a convenient analytical form for such a transformation. Consequently, deep learning architectures such as convolutional neural networks (CNNs) and recurrent neural networks (RNNs) have shown remarkable performance in BP estimation, by capturing the temporal and spatial nuances inherent in raw physiological signals [12]. Similarly, deep learning models have demonstrated significant potential in detecting AF from raw ECG signals. The authors of [13] developed a deep learning model employing a CNN to detect AF from single-lead ECG recordings. Their model demonstrated high accuracy, underscoring the potential of deep learning in the detection of arrhythmias. Similarly, [14, 15, 16] implemented deep learning approaches using CNNs on PPG signals for AF detection. These approaches yielded results that were competitive with ECG-based methods, thereby demonstrating the feasibility of utilizing PPG signals for AF detection. End-to-end deep learning frameworks that process raw ECG data to generate AF predictions have simplified and improved the accuracy of AF detection.
@@
Feature-Based Models. Besides working on raw data, another possibility to solve classification or regression tasks on PPG data is to establish machine learning models operating on clinically interpretable PPG features. These features include features based on pulse morphology, e.g. PPG pulse wave features such as the systolic peak, the diastolic peak or pulse duration and PPG derivative features (6,17), and irregularity features based on measures of randomness, variability and complexity in the inter-beat-intervals that can be determined from the PPG [17, 18]. These clinically interpretable PPG features are used for BP estimation (see below), AF detection (see below) and other questions related to the cardiovascular system, e.g. the assessment of arterial stiffness [19]. In addition, PPG signals show sometimes a strict periodic behaviour or, in general, a quasi-periodic behaviour. This motivates the use of features from Fourier- [20], Wavelet- [21] or Hilbert-Huang- [22] Transformations for training models. Such models are defined for BP estimation and AF detection. In related work, features from Fourier-Transformation are used to define models for the detection of aneurysms [23] or stenoses [24]. As PPG signals are in general not periodic, but quasiperiodic, one might expect better results with Wavelet- or Hilbert-Huang-transformations.
@@
Image-Based Models. Image-based models, such as those using the Continuous Wavelet Transform (CWT), convert physiological signals into visual representations, enabling deep learning to analyze, classify, and estimate physiological parameters. The CWT is a powerful tool for analyzing localized variations of power within a time series signal. Unlike the Fourier Transform, which provides a global frequency representation, CWT can provide a time-frequency representation that preserves the temporal localization of features. CWT based scalograms have already been used for PPG signal transformation to classify BP (Normal, Prehypertension, Stage 1 hypertension and Stage 2 hypertension) [25, 26], estimate heart rate variability (HRV) and signal quality [27] as well as to detect atrial fibrillation [15].
@@ 3.1 Datasets
Age (mean ± SD) 58.98 ± 15.03 58.89 ± 15.07
Sex (M%) 57.69 57.91
SBP (mmHg, mean ± SD) 115.48 ± 18.92 115.47 ± 18.91
DBP (mmHg, mean ± SD) 62.92 ± 12.08 62.93 ± 12.06
VitalDB dataset. The VitalDB dataset includes ECG, PPG, and invasive arterial blood pressure (ABP) signals collected using patient monitors from surgical patients [30]. Wang et al. (26) published the pre-processed VitalDB dataset as a subset of the PulseDB dataset, from which we use the PPG signals of 10s length with sampling frequency of 125 Hz, and reference systolic BP (SBP) and diastolic BP (DBP) values derived from the ABP signals. The dataset supports both calibration-based (in which train and test datasets share subjects) and calibration-free (in which train and test sets do not share subjects) testing approaches, which are essential for assessing the generalizability of BP estimation models. We refer to the first scenario as VitalDB ‘Calib’ and to the second scenario as VitalDB ‘CalibFree’. To ensure comparability with literature results, we keep the original test sets intact but split the original training sets into training, validation and calibration sets, where the latter is not considered in this study, mimicking the way the respective test sets were created, i.e., defining validation sets with/without patient overlap for VitalDB ‘Calib’/’CalibFree’. In Table LABEL:table_1, we summarize the two considered subsets, where one sample corresponds to a segment of 10s length.
Table 2: Characteristics of the DeepBeat subsets used for AF classification.
@@
DeepBeat dataset. For the AF classification task, we leveraged the DeepBeat dataset [29]. The dataset comprises more than 500,000 25-second, 32 Hz PPG segments, henceforth referred to as samples, from 175 individuals (108 with AF, 67 without). PPG signals were collected using a wrist-based PPG wearable device from cohorts of participants before cardioversion, patients undergoing an exercise stress test, and during daily life [29]. In the original publication [29], due to uneven AF/non-AF distribution across splits, performance metrics were overestimated, and the strong test scores did not reflect equivalent success on the validation and training sets. To address these issues, we implemented a new data split. We ensured no overlap between sets by redistributing subjects, thereby eliminating the redundancy present in the original dataset (Table LABEL:table_2). This revised split maintains an equal AF/non-AF ratio across the training, validation and test sets, providing a more reliable representation and enhancing the robustness of our model evaluation. (for more details please refer to the Supplementary Material).
@@ 3.2 Performance Evaluation and Metrics
Blood Pressure Estimation. We evaluate BP predictions using the Mean Absolute Error (MAE) and Mean Absolute Scaled Error (MASE), comparing model outputs against a baseline that predicts the training set median. Metrics are reported separately for systolic and diastolic pressures.
@@
  • Grade B: Errors between 5 and 6 mmHg and equal to 6 mmHg
  • Grade C: Errors between 6 and 7 mmHg and equal to 7 mmHg
@@
Atrial Fibrillation Detection. AF detection performance is assessed using standard metrics—sensitivity, specificity, receiver operator characteristic (ROC) area under curve (AUC), F1 score, and Matthews correlation coefficient (MCC). Classification thresholds are adjusted to meet desired sensitivity/specificity criteria, with a default threshold of 0.5 for the F1 score and optimized thresholds for other metrics. We explore two different threshold choices by fixing the threshold such that either sensitivity or specificity exceeds 0.8.

### v2

@@ Abstract
Background: Photoplethysmography (PPG) is a non-invasive physiological sensing method used in many clinical applications, increasingly supported by machine learning. However, systematic comparisons of input representations and models remain limited. Methods: We address this gap in the research landscape by a comprehensive benchmarking study covering three kinds of input representations, interpretable features, image representations and raw waveforms, across prototypical regression and classification use cases: blood pressure (BP) estimation and atrial fibrillation (AF) prediction. Results: In both cases, the best results are achieved by deep neural networks operating on raw time series as input representations. Within this model class, the strongest performance is observed for deeper convolutional neural networks (CNNs). However, depending on the task, smaller or lower-capacity CNNs can also achieve competitive performance, as confirmed by Bland–Altman analyses and statistical significance analyses based on bootstrapping. Conclusions: By providing a controlled, like-for-like comparison across signal, feature, and image-based representations, this study offers practical guidance for selecting robust machine-learning approaches for real-world PPG applications.
Keywords: Photoplethysmography, Machine Learning, Deep Neural Networks, Blood Pressure Estimation, Atrial Fibrillation Detection.
@@ 1 Introduction
PPG is one of the most commonly used wearable sensing techniques. It consists of projecting light onto the skin and measuring the amount of light that is reflected back or transmitted through the underlying tissues. Its simplicity, non-invasive nature, and ability to deliver multiple physiological parameters make it particularly attractive [8, 7]. As a result, PPG has been integrated into a range of clinical devices, such as pulse oximeters, as well as consumer wearable devices, including smartwatches.
The PPG signal measures the fluctuations in blood volume in the skin’s microvascular bed, which occur with each heartbeat. Figure 1 shows an exemplary PPG signal: the shape of the pulse wave contains information relating to the heart and vasculature, including BP, and the inter-beat intervals are related to heart rhythm [63]. The time delay between the electrical activation of the heart and the arrival of the corresponding pulse wave at a peripheral site where PPG is measured has been used to predict BP. There are generally two approaches to analysing PPG signals [8]: (i) using signal processing to extract features relating to pulse wave shape or inter-beat-intervals; and (ii) using deep learning techniques to analyse PPG signals or their image-based representations. Over the past few years, deep learning techniques have become widely used [17, 19, 50], as they have often shown superior performance to feature-based approaches.
@@
In the present study, we investigate two widely considered clinical applications for PPG analysis: AF classification as a prototypical classification task and cuffless BP estimation as a prototypical regression task. AF is the most common sustained cardiac arrhythmia and confers a five-fold increase in stroke risk [51]. AF is characterised by irregular and often very rapid heart rhythm. PPG provides an attractive approach to identifying AF because it is widely used in consumer wearables, and because it can detect each heartbeat, it can provide measures of the heart rhythm. During an AF episode, the time intervals between heartbeats are irregular. BP is one of the most widely used physiological measurements. It is a key marker of cardiovascular health; a valuable predictor of cardiovascular events; and is essential for the selection and monitoring of antihypertensive (BP lowering) treatments [23]. PPG-based BP estimation provides a potential approach to monitor BP unobtrusively in daily life. Accordingly, numerous studies have explored non-invasive and cuffless BP monitoring using PPG and related physiological signals, spanning feature-based methods as well as deep learning approaches [19, 50]. The rationale behind selecting two prototypical, but very different use cases, is to identify general patterns that could guide practitioners in the field even beyond the two investigated use cases.
While many prediction models have been put forward for the two considered prediction tasks at hand, benchmarking results are typically presented within a set of prediction models operating on a single kind of input modality, or a comparison is carried out against previously reported results from the literature. However, the latter relies on matching the experimental setup as closely as possible, where deviations from this setup severely limits the comparability of the results. The current lack of directly comparable benchmarking is an important gap in the research landscape, which we aim to address with this submission. This lack of controlled, directly comparable benchmarking across input representations and model families motivates the present study.
In this work, we address the following research questions: How do state-of-the-art machine learning models operating on different inputs representations compare? Are there universal patterns in terms of best-performing input representations or model architectures across different prototypical classification and regression use cases? As our main technical contribution, we put forward a like-for-like benchmarking of a comprehensive set of state-of-the-art algorithms covering three kinds of input representations on two large-scale datasets for a prototypical classification task (AF classification) and two different variants of a prototypical regression task (BP regression), using identical data splits and evaluation protocols to enable direct comparison across model families and to identify robust performance patterns that generalize beyond a single model or dataset.
@@ 2 Related Work
Overview of Learning Paradigms for PPG Analysis. To clarify the structure of this section, we organize prior work primarily by the input representation used for PPG analysis rather than by the learning algorithm. Specifically, we distinguish between models operating on raw time-series signals, engineered or clinically interpretable features, and image-based representations. While similar architectures (e.g., CNNs or RNNs) may appear across categories, the defining criterion is the form of the input representation. Typically, all approaches for clinical prediction tasks based on PPG data rely on a combination of signal processing and machine learning, where the precise focus varies considerably across different approaches. At one end of the spectrum, approaches relying on clinically interpretable features focus heavily on signal processing to extract meaningful features, and typically use comparably simple classification/regression models to perform the prediction. At the other end of the spectrum, deep learning methods using raw time series as input with as little signal processing as possible, rely on complex prediction models to extract and process meaningful features by themselves. In between, there are classifiers based on image-representations, that leverage signal processing tools to turn raw waveforms into image representations, and then most commonly also rely on deep learning models to perform the prediction.
@@
Raw Time Series as Input Representations. Recent advancements in deep learning have significantly impacted healthcare by enabling complicated analysis of raw physiological data. These models are particularly effective in estimating BP and detecting AF, among other applications. The key advantage of deep learning is its ability to recognize complicated patterns directly from raw data such as electrocardiogram (ECG) and PPG signals, eliminating the need for extensive manual feature development [77, 26, 68]. They provide an effective means to learn the complex, nonlinear underlying relationship between PPG signals and various physiological parameters, without the need to define a convenient analytical form for such a transformation. Consequently, deep learning architectures such as convolutional neural networks (CNNs) and recurrent neural networks (RNNs) have shown remarkable performance in BP estimation, by capturing the temporal and spatial nuances inherent in raw physiological signals [35]. Very recent work also demonstrated strong performance of hybrid CNN–BiLSTM architectures with attention on large datasets [47].
Similarly, deep learning models have demonstrated significant potential in detecting AF from raw ECG signals. The authors of [26] developed a deep learning model employing a CNN to detect AF from single-lead ECG recordings. Their model demonstrated high accuracy, underscoring the potential of deep learning in the detection of arrhythmias. Similarly, [37, 12, 2] implemented deep learning approaches using CNNs on PPG signals for AF detection. These approaches yielded results that were competitive with ECG-based methods, thereby demonstrating the feasibility of utilizing PPG signals for AF detection. End-to-end deep learning frameworks that process raw ECG data to generate AF predictions have simplified and improved the accuracy of AF detection.
Beyond PPG-specific investigations, several prior works [33, 14] have put forward benchmarks on a diverse set of time series classification tasks, establishing, for example, InceptionTime and MiniRocket as robust baselines models, which encourage their use in PPG-based prediction tasks.
More recently, transformer-based architectures have been explored for PPG and other physiological time-series analysis. Several studies have reported competitive performance of Transformer or hybrid CNN–Transformer models for time series forecasting [75, 41]. However, their effectiveness relative to convolutional architectures remains task- and data-dependent [20].
@@
Feature-Based Input Representations. Besides working on raw data, another possibility to solve classification or regression tasks on PPG data is to establish machine learning models operating on clinically interpretable PPG features. This includes features based on pulse morphology, e.g., PPG pulse wave features such as the systolic peak, the diastolic peak or pulse duration, and PPG derivative features [19, 46], and irregularity features based on measures of randomness, variability and complexity in the inter-beat-intervals that can be determined from the PPG [54, 18]. These clinically interpretable PPG features are used for BP estimation (see below), AF detection (see below), and other questions related to the cardiovascular system, e.g. the assessment of arterial stiffness [21]. In addition, PPG signals sometimes show a strict periodic behaviour or, in general, a quasi-periodic behaviour. This motivates the use of features from Fourier- [6], Wavelet- [45] or Hilbert-Huang- [31] Transformations for training models. Such models are defined for BP estimation and AF detection. In related work, features from Fourier-Transformation are used to define models for the detection of aneurysms [25] or stenoses [44]. As PPG signals are in general not periodic, but quasiperiodic, one might expect better results with Wavelet- or Hilbert-Huang-transformations.
Image-Based Input Representations. Image-based models, such as those using the Continuous Wavelet Transform (CWT) or short-time Fourier transform (STFT), convert physiological signals into visual representations, enabling deep learning to analyze, classify, and estimate physiological parameters.
@@
The CWT is a powerful tool for analyzing localized variations of power within a time series signal. Unlike the Fourier Transform, which provides a global frequency representation, CWT can provide a time-frequency representation that preserves the temporal localization of features. CWT-based scalograms have been used for PPG signal transformation to classify BP (Normal, Prehypertension, Stage 1 hypertension, and Stage 2 hypertension) [1, 76], estimate heart rate variability (HRV) and signal quality [49] as well as to detect AF [12]. CWT’s scale-adaptive time–frequency resolution suits PPG’s non-stationary, multi-scale nature, and its rich scalograms representation consistently delivers strong performance across various tasks.
The STFT-based spectrogram representation has been used in several PPG-based predictions using DL [11, 11]. In this approach, the PPG signals were transformed into fixed-resolution time–frequency spectrograms using the STFT and processed using a ResNet architecture. This approach, Spectrogram-ResNet, makes it possible to use the image-based convolutional neural networks by using a simpler time–frequency representation compared to the CWT process [11]. The standard STFT approach was also used in [11] for signal quality estimation.
Several alternatives have been explored for image-based PPG representations [12, 76]. These include direct conversion of the 1D PPG waveform into a structured 2D representation (e.g. multi-channel raw-PPG derivatives obtained by stacking padded PPG cycles (or beats)) which has also been applied for BP estimation [60], Gramian Angular Fields (GASF/GADF) [73] to encode temporal correlations, Visibility Graphs [72] providing a topological view of waveform structure, and Symmetric Projection Attractor Reconstruction (SPAR) attractor images which encapsulate the morphology and variability of the signal [3, 43]. The optimal choice depends on the downstream task, as each method has distinct strengths and limitations.
@@ 3.1 Datasets
VitalDB dataset. The VitalDB dataset includes ECG, PPG, and invasive arterial blood pressure (ABP) signals collected using patient monitors from surgical patients [40]. Wang et al. [71] published the pre-processed VitalDB dataset as a subset of the PulseDB dataset, from which we use the PPG signals of 10s length with a sampling frequency of 125 Hz, and reference systolic BP (SBP) and diastolic BP (DBP) values derived from the ABP signals. We used the VitalDB subset instead of the entire PulseDB dataset as it reduced the dataset size and was shown to lead to good generalization to out-of-distribution data [48].
The dataset supports both calibration-based and calibration-free evaluation protocols for assessing BP model generalizability, where calibration denotes subject-specific adaptation using reference BP measurements without retraining or fine-tuning the model, through samples from the same subject in the training and test sets following the PulseDB convention [71]. This differs from the calibration definitions used in some prior cuffless BP studies (e.g., [34]), where calibration may involve explicit model adaptation or fine-tuning. The calibration-free scenario ensures no patient overlap between training and test set and therefore assesses the generalization to unseen patients.
We refer to the first scenario as VitalDB ‘Calib’ and to the second scenario as VitalDB ‘CalibFree’. To ensure comparability with literature results, we keep the original test sets intact but split the original training sets into training, validation, and calibration sets, where the latter is not considered in this study, mimicking the way the respective test sets were created, i.e., defining validation sets with/without patient overlap for VitalDB ‘Calib’/’CalibFree’. In Table LABEL:table_1, we summarize the two considered subsets, where one sample corresponds to a segment of 10s length.
@@
Age (years, mean ± SD) 58.9 ± 15.0 58.8 ± 15.0
Sex (M%) 57.6 57.9
SBP (mmHg, mean ± SD) 115.4 ± 18.9 115.4 ± 18.9
DBP (mmHg, mean ± SD) 62.9 ± 12.0 62.9 ± 12.0
DeepBeat dataset. For the AF classification task, we leverage the DeepBeat dataset [67]. The dataset comprises more than 500,000 25-second, 32 Hz PPG segments, henceforth referred to as samples, from 175 individuals (108 with AF, 67 without). PPG signals were collected using a wrist-based PPG wearable device from cohorts of participants before cardioversion, patients undergoing an exercise stress test, and during daily life [67]. In the original publication [67], due to an uneven AF/non-AF distribution across splits, performance metrics were overestimated, and the strong test scores did not reflect equivalent success on the validation and training sets. To address these issues, we implement a new data split. We ensure no overlap between sets by redistributing subjects, thereby eliminating the redundancy present in the original dataset (Table LABEL:table_2). This revised split maintains an equal AF/non-AF ratio across the training, validation, and test sets, thus providing a more reliable representation and enhancing the robustness of our model evaluation. (For more details, please refer to the Supplementary Material.)
Table 2: Characteristics of the DeepBeat subsets used for AF classification. “Data Ratio” indicates the proportion of the full dataset assigned to each subset, and “AF Ratio” denotes the fraction of samples within that subset labeled as AF.
@@ 3.2 Performance Evaluation and Metrics
BP Estimation. We evaluate BP predictions using the Mean Absolute Error (MAE) and Mean Absolute Scaled Error (MASE), comparing model outputs against a baseline that predicts the training set median. Metrics are reported separately for systolic and diastolic pressures.
Furthermore, to determine the agreement between predictions and reference values in more detail, Bland-Altman analyses were also performed [5]. The bias was determined as the mean difference between predictions and reference values, as well as the corresponding limits of agreement (LoA). The standard deviation of these differences was multiplied by 1.96 to determine the LoA. While MAE roughly indicates the size of errors, bias directly indicates systematic errors, while the LoA gives insight into the variability of individual prediction errors.
@@
  • Grade B: Errors >> 5 mmHg and $\leq$ 6 mmHg
  • Grade C: Errors >> 6 mmHg and $\leq$ 7 mmHg
@@
We acknowledge that there are other BP estimation standards such as AAMI [64] and BHS [52]. In this manuscript, however, we restrict ourselves to the IEEE standard and defer a detailed comparison according to different BP estimation standards to future work.
AF Detection. AF detection performance is assessed using standard metrics: sensitivity, specificity, receiver operator characteristic (ROC) area under the curve (AUC), F1 score, and Matthews correlation coefficient (MCC). Classification thresholds are adjusted to meet desired sensitivity/specificity criteria, with a default threshold of 0.5 for the F1 score and optimized thresholds for other metrics. We explore two different threshold choices by fixing the threshold such that either sensitivity or specificity exceeds 0.8.
Statistical significance. For every experimental setting, the best model, defined as the model with the lowest MAE for the regression tasks or the highest AUC for the classification tasks, was taken as the reference model. We then computed the performance difference between the reference model and every other model by performing 1,000 bootstraps. A performance difference whose 95% confidence interval did not contain zero indicated a model to be statistically significantly worse. Conversely, if zero was in the interval, then the method indicated that the respective model was not statistically significantly worse than the reference, which was indicated in the result tables accordingly.